OpenStudy (anonymous):

A square patio has an area of 50 square feet. To the nearest tenth of a foot, what is the length of the patio? A. 7.5 ft B. 10.0 ft C. 12.5 ft D. 7.1 ft

jimthompson5910 (jim_thompson5910):

say we have a square with these 4 equal sides (all equal to some unknown number s) 

jimthompson5910 (jim_thompson5910):

the area of this square is \[\Large A = s*s\] \[\Large A = s^2\]

jimthompson5910 (jim_thompson5910):

we are told that the area of this particular square is 50 square feet, so A = 50 \[\Large A = s^2\] \[\Large 50 = s^2\] how do we isolate s?